OpenStudy (anonymous):

Okay, im seriously, NOT SMART :( Subtracting Binomials and monomials (6x-1)-(2x+3)

OpenStudy (anonymous):

6x-1)/5 - 2x/3 =3 you need to multiply both sides by 15 3(6x-1) - 5*2x = 45 18x -3 - 10x = 45 8x - 3 = 45 8x = 48 x = 48/8 = 6 now to check this is right, just put x=6 into the left side of your equation and see if it equals 3

OpenStudy (anonymous):

Im still confused :(

OpenStudy (anonymous):

I dont know any other way to explain it, but the answer is 3.